please help me solve this : !6 ^ 3/4
write an equivalent expression using radical notation.

Answer by mananth(16949)

16 ^ 3/4
..
=(2^4)^(3/4)
=2^(4*3/4)
=2^3
...
m.ananth@hotmail.ca